RawlsK29:

Need help on this problem; 4x + 8x = 9 - 6x I keep getting x = 2 but the answer book says its x = 1/2

RawlsK29:

@Vocaloid @Ultrilliam

Ultrilliam:

Can I see your steps?

RawlsK29:

Yes. 4x + 8x = 9 - 6x 12x = 9 - 6x +6x +6x 18x = 9 I am now confused about what i did

RawlsK29:

I am supposed to divide both sides by 18 right?

Ultrilliam:

Yes, sorry, was working it out.

RawlsK29:

Its fine, soo... omg im an idiot. Okay i see how it is 1/2 because 18/18 is 1 and 9/18 is 2

Ultrilliam:

There ya go =P

RawlsK29:

Thank you!(:

Ultrilliam:

Yw ^_^